OpenStudy (anonymous):

"2.5D motion" An airplane must drop supplies on a target, vertical distance = 400m, horizontal distance = 3500m, initial velocity = 250 m/s East. If the wind blows toward the South at 50 m/s, find the horizontal displacement while the package is in the air.

OpenStudy (anonymous):



OpenStudy (anonymous):

So first, I solved for the time for the package to hit the ground. dy = 0.5at^2 400 = 0.5(-9.8)t^2 t=9.04 seconds

OpenStudy (anonymous):

Then, the time the pilot needs to wait to drop supplies... dx = vit 3500 = 250t t=14 sec 14sec-9.04sec = 4.96sec before release

OpenStudy (raffle_snaffle):

When you solve for displacement I would think you need to take the two velocities and subtract them from each other..

OpenStudy (anonymous):

Then, the final velocity of the package: Vfy: vf^2 = vi^2 + 2ad vf^2 = 0^2 + 2ad vf^2 = 2(-9.8)(400) vf = 88.6m/s downward Vfx: vf=vi=250 m/s East Resolve: sqrt(88.6^2+250^2) = 265m/s Theta=tan^-1(88.6/250) = 19.5 degrees 265 m/s @ 19.5 degrees below horizontal

OpenStudy (anonymous):

So, If the wind blows toward the South at 50 m/s, find the horizontal displacement while the package is in the air.

OpenStudy (raffle_snaffle):

do you have the answers? I would like to try and solve this one as well on my own time.

OpenStudy (anonymous):

In the first equation you assumed that the initial velocity in the "y" direction was 0m/s, but doesn't the package has an initial velocity in the "y" axis? That velocity would be the 50m/s of the wind blowing south.

OpenStudy (anonymous):

The package has 0m/s in the downward direction to begin with, but the South is part D in the problem. South and y-velocity aren't the same because it's a different dimension. The solution in the notes: 3500-1240=2260m East d South: d=vt d=50(9.04) d=452 South Theta= tan^-1(452/2260) R=sqrt(452^2+2260^2) = 2305m @11.3 degrees South of East I just don't know where the 1240 came from.

OpenStudy (anonymous):

and why use 9.04 sec?

OpenStudy (raffle_snaffle):

Is that question at me or fg?

OpenStudy (anonymous):

Anyone lol. This just bothers me o-o

OpenStudy (raffle_snaffle):

I will think it through when I have my physics book handy. lol

OpenStudy (raffle_snaffle):

I forgot most of the 2D kinematic equations. Lol

OpenStudy (anonymous):